4. A plane is flying at 240 m s−1, on a bearing of 125º from due
north. Calculate its velocity component due south, and its
velocity component due east.
1
Expert's answer
2020-11-11T07:59:33-0500

Answer

Velocity (v) =240m/s

Angle"(\theta) =125-90=35°"

Components in East ="vcos\theta\\"

"240\times cos35°=196.5m/s"

Component in South

"=vsin\theta=240\times sin35°\\=137.6m/s"